# HalluVault: A Novel Logic Programming-aided Metamorphic Testing Framework for Detecting Fact-Conflicting Hallucinations in Large Language Models

### v1

@@ 4. Methodology
We design and implement HalluVault to address the aforementioned challenges, the workflow of which is illustrated in Figure 4. HalluVault is comprised of the following four modules, with each module to be detailed later.
Factual Knowledge Extraction (§4.1): Based on voluminous knowledge database dumps, HalluVault acquires fundamental information and factual triples of valid entities.
Logical Reasoning (§4.2): In this module, HalluVault leverages reasoning rules to generate sound and diverse facts as new ground truth knowledge.
@@ 4.1. Factual Knowledge Extraction
This step aims to extract fundamental facts from the input knowledge data into fact triples that can be utilized for logical reasoning. A fact triple is denoted as:
where and are entities in logic programming, and is a predicate. In this paper, we use the notation in Equation (missing) 5 to denote the static data and use the notation in Equation (missing) 1 to denote the action of querying about the predicate. Given a piece of knowledge in natural language, the key is to identify the , , and their .
Existing knowledge databases (Bollacker et al., 2007; Auer et al., 2007; Suchanek et al., 2007; Miller, 1995) not only encompass a vast array of documents and pages but also provide available structured data. Extracted from knowledge databases, the structured data would become an ideal resource for the construction and enrichment of factual knowledge. Thus, the genesis of our test case data is exclusively rooted in the entities and structured information sourced from current knowledge databases, ensuring a sophisticated and well-informed foundation for our testing framework. Basically, we follow the categorization of entities and relations used by WikiPedia (Auer et al., 2007) to perform the identification. Figure 5(a) shows the categories of the entities. Figure 5(b) shows the categories of the relations and some example fact triples. The detailed process is outlined in Algorithm 1. For specific entities and relations from various domains (Lines 2-3), we automatically query the database for constructing the ground truth triples (Lines 6-8), providing useful information for the following components.
@@ 4.2. Logical Reasoning
This step aims to derive additional, enriched information from previously extracted factual knowledge. HalluVault uses a logical programming-based processor to automatically generate new factual knowledge. This allows us to take one or more factual knowledge triples as input and generate a mutated triple as output with five types of inference rules.
To tackle the primary concern of generating FCH test cases with variability, we design five types of reasoning rules (Equation (missing) 4) prevalently adopted in several literature (Zhou et al., 2019; Ren and Leskovec, 2020; Liang et al., 2022; Tian et al., 2022; Abboud et al., 2020) in the context of knowledge reasoning. This provides sound strategies to prepare new facts for further test case generation. HalluVault will exhaustively apply all the rules to all their relevant fact triples to generate new knowledge. The definitions of the five types of rules are detailed as follows.
@@
Negation Reasoning Rule . Given a factual knowledge triple , then we can infer the new knowledge is not valid. indicates the negation of the relation .
@@
Rule#2: Symmetric Reasoning. In symmetric relationships, if the subject and object in a triple maintain coherence upon interchange, a new triple can be deduced in accordance with Definition 4.2.
@@
Rule#3: Inverse Reasoning. In an inverse relationship, the subject and object can be reversely linked through a variant of the original relation, as defined in Definition 4.3.
@@
Rule#4: Transitive Reasoning. In transitive relationships, if the object in one triple is the subject of the second triple, we can therefore derive a new triple following the Definition 4.4.
@@
Composite Reasoning Rules . Given multiple basic reasoning rules or predicates , we can use chain them up to form a new composite reasoning rule.
The process of applying these various rules to the ground truth triples extracted in the previous module can be referenced in Algorithm 2. An automatic rule generator could be designed at the first stage to generate prolog rules according to the relation type of the ground truth triples (as in Line 5-14). The corresponding query problems are also generated and mapped to the generated rules, which could be applied to the prolog query later. With the predetermined rules, we can be assisted with the prolog engine, asserting all the related triples and consulting the reasoning rules (Line 16-17). Then as stated in Line 20-24, the prepared query problems can be used to trigger the prolog inference after querying the prolog engine. By obtaining solutions from prolog, we can generate new knowledge triples based on the entities and their relationships provided.
@@ 4.3. Benchmark Construction
From the mutated triples, this module outlines our approach to constructing question-answer (Q&A) pairs and prompts to facilitate the automatic testing of FCH.
@@
Question Generation. To ensure effective and systematic test cases and prompt generation, we have adopted a method that utilizes entity relations mapping to predefined Q&A templates. In the construction of relation-based Q&A templates, one key aspect lies in aligning various types of relations with the corresponding question templates from the mutated triples, i.e., the predicate type in the triple. Different relation types possess unique characteristics and expressive requirements, leading to various predefined templates. As listed in Table 1, we map the relation types to question templates based on speech and the grammatical tense of the predicate, to guarantee comprehensive coverage. Beyond these universal templates, for hard-to-describe predicates, we employ customized templates to generate valid Q&A pairs. To enhance the construction of natural language formatted questions, we also leverage the LLM to refine the structure of Q&A pairs.
Another key aspect is regarding the automatic and reliable answer generation. We note that the answer to the corresponding question is readily attainable from the factual knowledge in the triple. Primarily, it is easy to determine whether the answer is true/false based on the mutated triples. Meanwhile, mutated templates with positive and negative semantics via the usage of synonyms or antonyms, which greatly enhance the question diversity, can be treated in a similar manner as the negation rule defined in Section 4.2. Specifically, if the answer to a question with original semantics is Yes/No, then for a question with mutated opposite semantics, the corresponding answer would naturally be the opposite, i.e., No/Yes. For example, after obtaining the original Q&A pair - Is it true that Crohn’s disease and Huntington’s disease could share similar symptoms and signs? - Yes., we can use some antonyms to mutate it into - Is it true that Crohn’s disease and Huntington’s disease have totally different symptoms and signs? - No.
Relation | Template | Example | |||||||
@@
Verb Phrase in Passive Voice |
Verb Phrase in Active Voice |
@@ 4.4. Response Evaluation
To achieve automated detection of factual consistency, our methodology first incorporates a parsing step that leverages advanced NLP techniques. This step is designed to extract essential semantic elements from each sentence within LLM outputs, assembling these elements into a coherent, semantic-aware structure. The foundational premise of our approach is predicated on evaluating the semantic similarity between these constructed structures, aiming to discern the degree of consistency in their underlying semantics. Subsequently, we propose the development of a set of similarity-based testing oracles. These oracles are instrumental in applying metamorphic testing principles, enabling us to systematically assess the consistency or inconsistency between LLM responses and the established ground truth. Our approach is structured around several critical steps, detailed as follows:
Step 1. Preliminary Screening. First, we eliminate scenarios in which the LLM declines to provide an answer, as indicated by the “answer” field of the LLM’s response (as described in Algorithm 3 Line 7-8). Most of these responses arise because the LLM lacks the relevant knowledge for the reasoning process. As these responses adhere to the LLM’s principle of honesty, we categorize them as normal responses.
Step 2. Response Parsing and Semantic Structure Construction. As stated in Algorithm 3 Line 10-11, for the remaining suspicious responses, we extract triples based on the statements contained in the reasoning process part of the LLM’s response. The extracted triples are used to construct a response semantic structure . In this structure, the entities (i.e., the subject and object) are depicted as nodes, while the relation between them is illustrated as an edge connecting these nodes. Concurrently, we construct a similar reference semantic structure using ground truth knowledge triples associated with the question.
Step 3. Similarity-based Metamorphic Testing and Oracles. We apply metamorphic relations to detect and evaluate potential errors in LLM responses, based on the relationships between inputs and outputs, rather than relying on traditional labeled data. In our context, metamorphic relations specifically refer to comparing the similarity between semantic structures generated by LLMs and the ground truth counterparts, to identify and classify hallucination answers from LLMs (as mentioned in Algorithm 3 Line 12-18).
@@
where and denote the set of edges extracted from and , and is a predefined threshold (to be detailed in Section 5.1). Intuitively, the similarity score is calculated as the proportion of identical edges shared between the two sets against the total number of unique edges in both sets. If the similarity score is smaller than the threshold, then an FCH is detected. Note that when determining the joint and union of sets and , we consider two edges as identical if their corresponding relations are identical or represented by synonymous words, and vice versa.
@@
where and denotes the set of nodes extracted from and , and is a predefined threshold (to be detailed in Section 5.1). Intuitively, the similarity score is calculated as the proportion of identical nodes shared between the two sets against the total number of unique nodes in both sets. If the similarity score is smaller than the threshold, then an FCH is detected. Note that when determining the joint and union of sets and , we consider two nodes as identical if their corresponding entities are identical or represented by synonymous words, and vice versa.
@@ 5. Evaluation
RQ1 (Effectiveness): How effective is HalluVault for identifying LLM FCH issues? This RQ studies the effectiveness of HalluVault in generating test cases and identifying LLM FCH issues.
RQ2 (Hallucination Categorization and Analysis): What is the categorization of LLM FCH issues? This RQ categorizes the FCH issues of various LLMs identified by HalluVault. We also provide case studies for some specific cases, including temporal-related hallucination responses, out-of-distribution-data knowledge-related hallucination responses, and refusal normal responses.
RQ3 (Ablation Study): Whether the four types of logic reasoning rules can identify LLM FCH issues independently? This RQ explores whether the logic reasoning rules of HalluVault can effectively identify LLM FCH issues separately.
@@ 5.1. Experimental Setup
Logic Reasoning Processor. For the logic reasoning module, we apply SWI-Prolog, an open-source advanced logical programming interpreter. To effectively prevent errors due to excessive stacked strings, and ensure the proper operation of the logical processor when inserting a large number of facts into Prolog, we employ a sampling method and extract a subset of entities to form a query.
@@
Models Under Test. To guarantee a reliable evaluation for RQ1 and RQ2, we evaluate six state-of-the-art large language models with HalluVault. Considering the diverse nature of LLMs, we select two distinct categories for in-depth analysis: the first category comprises API-accessible models with closed-source architecture including ChatGPT (gpt-3.5-turbo-0613) and GPT-4 (OpenAI, 2023), and the second category consists of open-source LLMs with deployability, including Llama2-7B-chat-hf, Llama2-70B-chat-hf (Touvron et al., 2023), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024).
@@
Response validation Threshold . To validate responses from LLMs as described in Section 4.4, we apply StanfordOpenIE (Angeli et al., 2015) for knowledge triple extraction from LLM responses and then use SentenceBERT (Reimers and Gurevych, 2019) to calculate the vector similarity of nodes and edges from the constructed semantic structures. Here we set the threshold to 0.8, considering knowledge triples as semantically equivalent if they exceed this threshold, and vice versa. To determine the threshold value, we sample 30 test cases and corresponding LLM responses from each of the nine knowledge domains listed in Figure 5(a). Through this analysis, we find that by setting the threshold values for both and at 0.8, our oracles can achieve 100% precision in detecting non-equivalent LLM answers as FCHs. In other words, all instances where an LLM’s answer have a semantic similarity score below 0.8 compared to the ground truth were correctly identified as FCH cases.
@@ 5.2. RQ1: Effectiveness
To reveal the effectiveness of HalluVault, we evaluate the statistics of test cases generated by HalluVault and then evaluate the capabilities of identifying LLM FCH issues with the generated test cases. To further assess the effectiveness of test cases for uncovering FCH issues in specific knowledge domains, we evaluate the performances of LLMs on test cases across various knowledge domains.
Effectiveness on Generating Q&A Test Cases. We apply HalluVault to generate a Q&A test benchmark, amounting to a comprehensive total of 7,200 test cases, designed to provide a broad and detailed evaluation of LLM FCH issues across specific knowledge domains.
Effectiveness across LLMs. We instruct LLMs under test utilizing Q&A pairs derived from HalluVault, and automatically label both hallucination and normal responses. The results are presented in Figure 6(a), illustrating the proportion of hallucination responses versus normal responses from LLMs under test.
Among all models, GPT-4 exhibits the best performance, demonstrating the lowest proportion of hallucinatory responses in the test cases generated by HalluVault, at only 24.7%, while ChatGPT falls slightly behind with 42.1%. Open-source LLMs including Llama2-7B-chat-hf and Mistral-7B-Instruct-v0.2 with fewer parameters perform worse, but their counterparts with larger parameters (i.e., Llama2-70B-chat-hf and Mixtral-8x7B-Instruct-v0.1) achieve higher normal response rates surpassing ChatGPT on HalluVault. This indicates that the test cases generated by HalluVault successfully trigger hallucination responses across various LLMs when confronted with questions requiring logical reasoning capabilities.
Effectiveness on Specific Domain Knowledge for Each LLM. To further explore the effectiveness of HalluVault in identifying FCH issues spanning various domains of LLMs, we compare hallucination response across nine specific domain knowledge. Figure 6(b) presents the generated heatmaps of the confusion matrices for the specific knowledge field hallucination response rate based on the testing results. It can be clearly observed that different models exhibit varying strengths and weaknesses across distinct knowledge domains.
An interesting finding is that, within the domains of natural sciences and mathematics, LLMs generally exhibit weaker performance. This is potentially because there are many astrophysical or mathematical entities and their interrelationships in generated test cases by HalluVault. To answer such questions, the LLM needs an extensive understanding of astrophysical knowledge and mathematical theory. Thus, we infer that this realm of knowledge is not well-covered in the training datasets of LLMs under test, thereby resulting in high hallucination rates. Such a disparity in knowledge is likely a significant factor in the observed underperformance of LLMs in these specific domains.

### v2

@@ 4. Methodology
We design and implement Drowzee to address the aforementioned challenges, the workflow of which is illustrated in Figure 4. Drowzee is comprised of the following four modules, with each module to be detailed later.
Factual Knowledge Extraction (§4.1): Based on voluminous knowledge database dumps, Drowzee acquires fundamental information and factual triples of valid entities.
Logical Reasoning (§4.2): In this module, Drowzee leverages reasoning rules to generate sound and diverse facts as new ground truth knowledge.
@@ 4.1. Factual Knowledge Extraction
This step aims to extract fundamental facts from the input knowledge data into fact triples that can be utilized for logical reasoning.
Existing knowledge databases (Bollacker et al., 2007; Auer et al., 2007; Suchanek et al., 2007; Miller, 1995) not only encompass a vast array of documents and pages but also provide available structured data. Extracted from knowledge databases, the structured data would become an ideal resource for the construction and enrichment of factual knowledge. Thus, the genesis of our test case data is exclusively rooted in the entities and structured information sourced from current knowledge databases, ensuring a sophisticated and well-informed foundation for our testing framework. Basically, we follow the categorization of entities and relations used by WikiPedia (Auer et al., 2007) to perform the identification. Figure 5(a) shows the categories of the entities. Figure 5(b) shows the categories of the relations and some example fact triples.
The detailed process is outlined in Algorithm 1. As defined in the previous Equation 2, we extract the facts in the structure of three-element predicates, i.e., , where “” (stands for ) and “” (stands for ) are entities, and “” is the name of the predicate. The facts extraction is done on a per-category basis, implementing a divide-and-conquer strategy, which efficiently integrates all the facts ranging from all the categories. As shown in Algorithm 1, for any given entity category and relation category, the function ExtractGroundFacts iterates through all possible entities and relations. For each combination (), it queries the database using the QueryDB function (Lines 3-6), which retrieves all three-element facts established with the specific predicate and the argument placed either in the subject or the object position.
@@ 4.2. Logical Reasoning
This step aims to derive additional, enriched information from previously extracted factual knowledge. Drowzee uses a logical programming-based processor to automatically generate new factual knowledge. This allows us to take one or more factual knowledge triples as input and generate a derived triple as output with five types of inference rules.
To tackle the primary concern of generating FCH test cases with variability, we design five types of reasoning rules (Equation (missing) 4) prevalently adopted in several literature (Zhou et al., 2019; Ren and Leskovec, 2020; Liang et al., 2022; Tian et al., 2022; Abboud et al., 2020) in the context of knowledge reasoning. This provides sound strategies to prepare new facts for further test case generation. Drowzee will exhaustively apply all the rules to all their relevant fact triples to generate new knowledge. The definitions of the five types of rules are detailed as follows.
@@
Negation Reasoning Rule . Given a factual knowledge triple , then we can derive the new triple is not valid. indicates the negation of the relation .
@@
Rule#2: Symmetric Reasoning. In symmetric relations, if the subject and object in a triple maintain coherence upon interchange, a new triple can be deduced in accordance with Definition 4.2.
@@
Rule#3: Inverse Reasoning. In an inverse relation, the subject and object can be reversely linked through a variant of the original relation, as defined in Definition 4.3.
@@
Rule#4: Transitive Reasoning. In transitive relations, if the object in one triple is the subject of the second triple, we can therefore derive a new triple following the Definition 4.4.
@@
Composite Reasoning Rules . Given multiple basic reasoning rules or predicates , we can chain them up to form a new composite reasoning rule.
The process of applying these various rules to the ground truth triples extracted in the previous module can be referenced in Algorithm 2. An automatic rule generator could be designed at the first stage to iterate its predicates and generate the derivation rules according to the relation category (as in Lines 3-4). The corresponding query problems are also generated and mapped to the generated rules, which could be applied to the Prolog query later. With the predetermined rules, we can be assisted with the Prolog engine, asserting all the related triples and consulting the reasoning rules, as outlined in Lines 5-6. We use to denote the query results of w.r.t the Prolog program . When contains no variables, it returns Boolean results indicating the presence of the fact; otherwise, it outputs all the possible instantiations of the variables. Then as stated in Lines 7-9, by obtaining solutions from Prolog, we can generate new knowledge triples based on the entities and their relations provided. For each instantiation that contains one subject “s” and one object “o”, we then compose them with the new predicate, which is taken as one derived fact. These derived facts are later used to generate test cases.
@@ 4.3. Benchmark Construction
From the derived triples, this module outlines our approach to constructing question-answer (Q&A) pairs and prompts to facilitate the automatic testing of FCH.
@@
Question Generation. To ensure effective and systematic test cases and prompt generation, we have adopted a method that utilizes entity relations mapping to predefined Q&A templates. In the construction of relation-based Q&A templates, one key aspect lies in aligning various types of relations with the corresponding question templates from the derived triples, i.e., the predicate type in the triple. Different relation types possess unique characteristics and expressive requirements, leading to various predefined templates. As listed in Table 1, we map the relation types to question templates based on speech and the grammatical tense of the predicate, to guarantee comprehensive coverage. Beyond these universal templates, for hard-to-describe predicates, we employ customized templates to generate valid Q&A pairs. To enhance the construction of natural language formatted questions, we also leverage the LLM to refine the structure of Q&A pairs.
Another key aspect is regarding the automatic and reliable answer generation. We note that the answer to the corresponding question is readily attainable from the factual knowledge in the triple. Primarily, it is easy to determine whether the answer is true/false based on the derived triples. Meanwhile, mutated templates with positive and negative semantics via the usage of synonyms or antonyms, which greatly enhance the question diversity, can be treated in a similar manner as the negation rule defined in Section 4.2. Specifically, if the answer to a question with original semantics is Yes/No, then for a question with mutated opposite semantics, the corresponding answer would naturally be the opposite, i.e., No/Yes. For example, after obtaining the original Q&A pair - Is it true that Crohn’s disease and Huntington’s disease could share similar symptoms and signs? - Yes., we can use some antonyms to mutate it into - Is it true that Crohn’s disease and Huntington’s disease have totally different symptoms and signs? - No.
Relation | Template Examples | ||||
@@ 4.4. Response Evaluation
To achieve automated detection of factual consistency, our methodology first incorporates a parsing step that leverages advanced NLP techniques. This step is designed to extract essential semantic elements from each sentence within LLM outputs, assembling these elements into a coherent, semantic-aware structure. The foundational premise of our approach is predicated on evaluating the semantic similarity between these constructed structures, aiming to discern the degree of consistency in their underlying semantics. Subsequently, we propose the development of a list of similarity-based testing oracles. These oracles are instrumental in applying metamorphic testing principles, enabling us to systematically assess the consistency or inconsistency between LLM responses and the established ground truth. Note that our focus is on the accuracy of ground truth facts rather than highly specialized or sequential content like mathematical proofs. Consequently, during evaluation, we emphasize whether the entities and relations in the response align with the ground truth, regardless of the order in which the facts are presented. Our approach is structured around several critical steps, detailed as follows:
Step 1. Preliminary Screening. First, we eliminate scenarios in which the LLM declines to provide an answer, as indicated by the “answer” field of the LLM’s response (as described in Algorithm 3 Lines 3-4). Most of these responses arise because the LLM lacks the relevant knowledge for the reasoning process. Since these responses adhere to the LLM’s principle of honesty, we classify them as correct and normal responses, denoted as in the algorithm.
Step 2. Response Parsing and Semantic Structure Construction. As stated in Algorithm 3 Lines 6-7, for the remaining suspicious responses, the ExtractTriple function is used to generate triples based on the statements contained in the reasoning process part of the LLM’s response. Then from the extracted triples (), the BuildGraph function can construct a semantic structure , where the (i.e., the subject and object) are represented as nodes, and the between them is illustrated as an edge connecting these nodes. Concurrently, the ground truth triples () associated with the question are used as input to construct a similar semantic structure .
Step 3. Similarity-based Metamorphic Testing and Oracles. We apply metamorphic relations to detect and evaluate potential errors in LLM responses, based on the relationships between inputs and outputs, rather than relying on traditional labeled data. In our context, metamorphic relations specifically refer to comparing the similarity between semantic structures generated by LLMs and the ground truth counterparts, to identify and classify hallucination answers from LLMs (as mentioned in Algorithm 3 Lines 8-17). Note that we provide four classifications: correct responses (denoted as ), hallucinations caused by error inference (), hallucinations caused by erroneous knowledge (), and hallucinations containing both issues ().
@@
where and denote the list of edges extracted from and , and is a predefined threshold (to be detailed in Section 5.1). Intuitively, the similarity score is calculated as the proportion of identical edges shared between the two lists against the total number of unique edges in both lists. If the score is smaller than the threshold, then an FCH is detected. Note that when determining the joint and union of lists and , we consider two edges as identical if their corresponding relations are identical or represented by synonymous words, and vice versa.
@@
where and denotes the list of nodes extracted from and , and is a predefined threshold (to be detailed in Section 5.1). Intuitively, the similarity score is calculated as the proportion of identical nodes shared between the two lists against the total number of unique nodes in both lists. If the score is smaller than the threshold, then an FCH is detected. Note that when determining the joint and union of lists and , we consider two nodes as identical if their corresponding entities are identical or represented by synonymous words, and vice versa.
@@ 5. Evaluation
RQ1 (Effectiveness): How effective is Drowzee for identifying LLM FCH issues? This RQ studies the effectiveness of Drowzee in generating test cases and identifying LLM FCH issues.
RQ2 (Hallucination Categorization and Analysis): What is the categorization of LLM FCH issues? This RQ categorizes the FCH issues of various LLMs identified by Drowzee. We also provide case studies for some specific cases, including temporal-related FCHs and out-of-distribution-data knowledge-related FCHs.
RQ3 (Comparison with Existing Works): How does Drowzee compare with existing approaches in detecting LLM FCH issues? This RQ investigates whether Drowzee outperforms existing testing benchmarks and methods in constructing test cases and identifying LLM FCH issues. We conduct a qualitative analysis as well as a small-scale quantitative analysis of the accuracy of current hallucination detection methods compared with Drowzee.
RQ4 (Ablation Study): Whether the four types of logic reasoning rules can identify LLM FCH issues independently? This RQ explores whether the logic reasoning rules of Drowzee can effectively identify LLM FCH issues separately.
@@ 5.1. Experimental Setup
Logic Reasoning Processor. For the logic reasoning module, we apply SWI-Prolog (Wielemaker et al., 2012), an open-source advanced logical programming interpreter. To effectively prevent errors due to excessive stacked strings, and ensure the proper operation of the logical processor when inserting a large number of facts into Prolog, we employ a sampling method and extract a subset of entities to form a query.
@@
Models Under Test. To guarantee a reliable evaluation for RQ1 and RQ2, we evaluate six state-of-the-art large language models with Drowzee. Considering the diverse nature of LLMs, we select two distinct categories for in-depth analysis: the first category comprises API-accessible models with closed-source architecture including ChatGPT (gpt-3.5-turbo-0613) and GPT-4 (OpenAI, 2023), and the second category consists of open-source LLMs with deployability, including Llama2-7B-chat-hf, Llama2-70B-chat-hf (Touvron et al., 2023), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024).
Model Configuration. We set the temperature parameter to 0 to achieve more stable and conservative model outputs, ensuring consistency in the content generated during LLM testing. Additionally, we set the top-p value to 0.9 and disable top-k (set to 0) to filter out low-probability tokens and select the most likely tokens, thereby improving the accuracy of the generated results. To further validate the consistency of the LLM responses, we conducted significance tests to validate the consistency of the LLM responses. Specifically, we randomly selected 40 test cases from each domain under each rule, resulting in a total of 1440 test cases. Using GPT-3.5-turbo as an example, each test case is run five times under the previously described configuration to interact with the LLM. We then use the Sentence-bert model (Reimers and Gurevych, 2019) to calculate the pairwise cosine similarity between the five LLM responses generated for each test case. The consistency of the LLM responses is evaluated using Friedman tests (ScienceDirect, 2024), a non-parametric statistical test commonly employed to detect differences in treatments across multiple test attempts. The results show no significant differences between the responses of different runs, with an approximate average p-value of 0.54. This confirms that the generated responses are consistent across multiple executions, justifying the use of a single run for evaluation purposes.
@@
Response Validation Threshold . To validate responses from LLMs as described in Section 4.4, we apply StanfordOpenIE (Angeli et al., 2015; Remy, 2020) for knowledge triple extraction from LLM responses and then use Phrase-BERT (Wang et al., 2021) to calculate the vector similarity of nodes and edges from the constructed semantic structures. We also utilize GPT-4 to extract triples for some complex responses that StanfordOpenIE cannot handle effectively. Here we set the threshold to 0.8, considering knowledge triples as semantically equivalent if they exceed this threshold, and vice versa. To determine the threshold value, we sample 30 test cases and corresponding LLM responses from each of the nine knowledge domains listed in Figure 5(a). Through this analysis, we find that by setting the threshold values for both and at 0.8, with the given 270 test cases that are correctly classified, we can estimate the true positives among all test cases through Laplace’s approach in the Sunrise problem (Laplace, 1951), resulting in 99.6% when detecting non-equivalent LLM answers as FCHs. In other words, all instances where an LLM’s answer has a semantic similarity score below 0.8 compared to the ground truth were correctly identified as FCH cases.
Consistency of Words. To ensure word consistency in experiments, we maintain several dictionaries. For symmetric relations, we have a dictionary that includes relations which retain their meaning when the subject and object are reversed. Additionally, we use synonym dictionaries provided by NLP libraries (e.g., WordNet (Miller, 1994)) along with our own set of synonyms tailored for specific cases when validating the LLM responses.
@@ 5.2. RQ1: Effectiveness
To reveal the effectiveness of Drowzee, we evaluate the statistics of test cases generated by Drowzee and then evaluate the capabilities of identifying LLM FCH issues with the generated cases. To further assess the effectiveness of test cases for uncovering FCH issues in specific knowledge domains, we evaluate the performances of LLMs on test cases across various knowledge domains.
Effectiveness on Generating Q&A Test Cases. We apply Drowzee to generate a Q&A test benchmark, amounting to a comprehensive total of 7,200 test cases, designed to provide a broad and detailed evaluation of LLM FCH issues across specific knowledge domains.
Effectiveness across LLMs. We instruct LLMs under test utilizing Q&A pairs derived from Drowzee and automatically label both hallucination and normal responses. Different LLMs might trigger different hallucinations on the same questions. The results are presented in Figure 6(a), illustrating the proportion of FCHs versus normal responses from LLMs under test.
Among all models, GPT-4 exhibits the best performance, demonstrating the lowest proportion of hallucinatory responses in the test cases generated by Drowzee, at only 24.7%, while ChatGPT falls slightly behind with 42.1%. Open-source LLMs including Llama2-7B-chat-hf and Mistral-7B-Instruct-v0.2 with fewer parameters perform worse, but their counterparts with larger parameters (i.e., Llama2-70B-chat-hf and Mixtral-8x7B-Instruct-v0.1) achieve higher normal response rates surpassing ChatGPT on Drowzee. This indicates that the test cases generated by Drowzee successfully trigger hallucination responses across various LLMs when confronted with questions requiring logical reasoning capabilities.
Effectiveness on Specific Domain Knowledge for Each LLM. To further explore the effectiveness of Drowzee in identifying FCH issues spanning various domains of LLMs, we compare hallucination response across nine specific domain knowledge. Figure 6(b) presents the generated heatmaps of the confusion matrices for hallucination response rate from the specific knowledge field based on the testing results. It can be clearly observed that different models exhibit varying strengths and weaknesses across distinct knowledge domains.
An interesting finding is that, within the domains of natural sciences and mathematics, LLMs generally exhibit weaker performance. This is potentially because there are many astrophysical or mathematical entities and their interrelationships in generated test cases by Drowzee. To answer such questions, the LLM needs an extensive understanding of astrophysical knowledge and mathematical theory. Thus, we infer that this realm of knowledge is not well-covered in the training datasets of LLMs under test, thereby resulting in high hallucination rates. Such a disparity in knowledge is likely a significant factor in the observed underperformance of LLMs in these specific domains.